OpenStudy (anonymous):

Someone please help. Salina currently has an account balance of $1,047.69. Her initial deposit on the account was $630 and it earned 3.9% simple interest. How long has Salina held the account?

OpenStudy (anonymous):

ok so if P is principal for one period she gets P*3.9% simple interest.

OpenStudy (anonymous):

formula is I = P r t, where r=% in decimal t=periods(years) P=Principal

OpenStudy (anonymous):

"Her initial deposit on the account was $630" P = $630

OpenStudy (otonogold):

A = \frac{P \times r \times t}{100} A = 1047.69 P = 630 r = 3.9 Put these values and solve for t...

OpenStudy (otonogold):

Soon as you set up your problem Multiply by 100 both sides

OpenStudy (otonogold):

104769=2457x 42 <-

OpenStudy (anonymous):

formula is A= P r t; not I above.